OpenStudy (anonymous):

A curve is traces by a point P(x,y) which moves such that its distance from the point A(1, -2) is twice its distance from the point B(4, -3). Determine the equation of the curve

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

If \(d_1\) is the distance between P and A, and \(d_2\) is the distance between P and B, then you have the following system of equations: \[\begin{cases}d_1=\sqrt{(x-1)^2+(y+2)^2}\\d_2=\sqrt{(x-4)^2+(y+3)^2}\\d_1=2d_2\end{cases}\] It looks like you must solve for \(y\) in terms of \(x\).

OpenStudy (anonymous):

Thank you but I am still kinda confused so to go further to find the equation you would?